# F-95000 Cergy-Pontoise, France

### v1

@@ 8 Concluding Remarks
Data analytics requires the collection and integration of data coming from different sources and stored in what is called a data warehouse that is a database operating under a specific type of schema called a star schema. As a consequence, more often than not, such integrated data contain inconsistencies and missing values. In this paper, we have seen an approach to efficiently compute consistent answers to analytic queries in data warehouses. Extending earlier work concerning the consistent query answering of standard, non-analytic queries in multi-table databases, we presented in this paper an approach to computing exact consistent answers of analytic queries over star schemas under two assumptions: (a) the selection condition in the query involves no keys and (b) the selection condition is independent. Our main contributions are: (a) a polynomial algorithm for computing the exact consistent answer to a usual projection-selection-join query over a star schema, and (b) showing that the exact consistent answer to an analytic query over a star schema can be computed also in time polynomial in the size of the data warehouse.
Our current work follows two main lines of research. First, the implementation of our approach is an important issue under consideration. We emphasize in this respect that since the functional dependencies in a star schema have a simple form, the generic algorithm $m\_Chase$ as presented here can be optimized. Moreover, as $m\_Chase(T)$ has to be ‘re-computed’ after each update, the question of incrementally implementing the changes in $m\_Chase(T)$ is crucial regarding efficiency. We are currently investigating these issues.
@@ Appendix A Proof of Proposition 6
Indeed, let $X\to A$ in $FD$, $t_{\varphi}(\sigma)$ and $t_{\varphi}(\sigma^{\prime})$ in $T_{\varphi}$ such that $t_{\varphi}(\sigma).X=t_{\varphi}(\sigma^{\prime}).X$. If $\sigma$ and $\sigma^{\prime}$ are defined over $A$, then, by construction of $T_{\varphi}$, we have $t_{\varphi}(\sigma).A=t_{\varphi}(\sigma^{\prime}).A$, in which case $m\_Chase$ does not change $T_{\varphi}$. The case whereby $\sigma$ is defined over $A$ and $\sigma^{\prime}$ is not defined over $A$ is not possible in $m\_Chase(T)$, and thus it is not possible that $t_{\varphi}(\sigma)$ is defined over $A$ while $t_{\varphi}(\sigma^{\prime})$ is not. Therefore, we have $m\_Chase(T_{\varphi})=T_{\varphi}$, and this implies that $T_{\varphi}\models FD$, because no conflicts can occur in $T_{\varphi}$.
@@
Recalling from LauS25 that, when $FD$ is acyclic, we also have $m\_Chase({\sf Cons}(\mathcal{T}))={\sf Cons}(\mathcal{T})$ and ${\sf Cons}(\mathcal{T})\models FD$. We prove by induction that, at each step $k$ of the computation of $m\_Chase(R_{\varphi})$ the obtained m-table $\Sigma^{k}$ is such that $(1)$ $\Sigma^{k}$ contains no conflict and $(2)$ $\tau(\Sigma^{k})=\tau(R_{\varphi})$.
@@
-- If $q$ and $q^{\prime}$ are both in $T_{\varphi}$ (respectively both in ${\sf Cons}(\mathcal{T})$): as these sets are ‘closed’ under m_Chase, $t$ and $t^{\prime}$ are both in $T_{\varphi}$ (respectively both in ${\sf Cons}(\mathcal{T})$), and so $t^{\prime}_{A}$ is in $T_{\varphi}$ (respectively in ${\sf Cons}(\mathcal{T})$), because $t^{\prime}_{A}$ is obtained through the m-Chase process.
-- If $q$ is in ${\sf Cons}(\mathcal{T})$ and $q^{\prime}$ is in $T_{\varphi}\cap{\sf Confl}(\mathcal{T})$: in this case $xa$ is in ${\sf Cons}(\mathcal{T})$ (because $xa\sqsubseteq q$). If $\sigma^{\prime}$ is the m-tuple in $m\_Chase(T)$ such that $q^{\prime}=t_{\varphi}(\sigma^{\prime})$, $XA\subseteq sch(\sigma^{\prime})$ and $xa$ is in ${\sf tuples}(\sigma^{\prime}(XA))$. Hence $t^{\prime}_{A}\sqsubseteq t_{\varphi}(\sigma^{\prime})$, and thus $t^{\prime}_{A}$ is in ${\sf tuples}(R_{\varphi})$.
-- If $t$ is in $T_{\varphi}\cap{\sf Confl}(\mathcal{T})$ and $t^{\prime}$ in ${\sf Cons}(\mathcal{T})$: in this case, assume first that $xa$ is in ${\sf Cons}(\mathcal{T})$. Then, $t^{\prime}_{A}$ is in ${\sf Cons}(\mathcal{T})$, because for every $Y\to B$ in $FD$ other than $X\to A$, if $YB\subseteq sch(t^{\prime}_{A})$ then $YB\subseteq sch(t^{\prime})$.
Assuming now that $xa$ is in ${\sf Confl}(\mathcal{T})$, we have $a=\varphi(x)$ and so, for every $Y\to B$ such that $YB\subseteq sch(t^{\prime}_{A})$, $t^{\prime}_{A}.B=\varphi(t^{\prime}_{A}.Y)$ (because as $t^{\prime}$ is in ${\sf Cons}(\mathcal{T})$, we also have that for every $Y\to B$ such that $YB\subseteq sch(t^{\prime})$, $t^{\prime}.B=\varphi(t^{\prime}.Y)$). Hence, there exists $\sigma^{\prime}$ in $m\_Chase(T)$ such that $t^{\prime}_{A}\sqsubseteq t_{\varphi}(\sigma^{\prime})$. Therefore $t^{\prime}_{A}$ is in $\tau(R_{\varphi})$.
@@
To this end, let $S$ be such that ${\sf True}(\mathcal{R}_{\varphi})\subseteq{\sf True}(\mathcal{S})\subseteq{\sf True}(\mathcal{T})$ and $S\models FD$, and we consider a tuple $q$ in ${\sf True}(\mathcal{S})\setminus{\sf True}(\mathcal{R}_{\varphi})$. Then, as $q$ is in ${\sf True}(\mathcal{T})$, there exist $\sigma$ in $m\_Chase(T)$ and $t$ in ${\sf tuples}(\sigma)$ such that $q\sqsubseteq t$. Since ${\sf Cons}(\mathcal{T})\subseteq{\sf True}(\mathcal{R}_{\varphi})$, it follows that $q$ is in ${\sf Confl}(\mathcal{T})$, implying that there exists $X\to A$ in $FD$ such that $XA\subseteq sch(q)$ and ${\sf True}(\mathcal{T})$ contains $xa^{\prime}$ such that $x=q.X$ and $a^{\prime}\neq q.A$. By construction of $R_{\varphi}$, ${\sf True}(\mathcal{R}_{\varphi})$ contains a tuple $t^{\prime}$ from ${\sf tuples}(\sigma)$, and so, we have $t.X=q.X=t^{\prime}.X$ (because for every $X\to A\in FD$, $|{\sf tuples}(\sigma(X))|=1$) and $t.A\neq t^{\prime}.A$. As $t.A=q.A$ and $t^{\prime}$ is in ${\sf True}(\mathcal{S})$, this implies that $S\not\models FD$, which is a contradiction. This part of the proof is therefore complete.
@@
Appendix B Proof of Proposition 8
Proposition 8  Let $T$ be a star-table over universe $U$, and $Q:{\tt select}$ $X$ from $T$ where $\Gamma$ a query, such that $\mathbf{K}\cap sch(\Gamma)=\emptyset$ and $\Gamma$ is an independent selection condition. Then ${\sf C\_ans}(Q)$ is the set of all tuples $x$ over $X$ for which there exists $\sigma$ in $m\_Chase(T)$ such that

### v2

@@ 8 Concluding Remarks
Data analytics requires the collection and integration of data coming from different sources and stored in a database operating under a specific type of schema called a ‘star schema’. Such a database is referred to in the literature as a data warehouse inmon96 . As a consequence, more often than not, such integrated data contain inconsistencies and missing values. In this paper, we have seen an approach to efficiently compute consistent answers to analytic queries in data warehouses. Extending earlier work concerning the consistent query answering of standard, non-analytic queries in multi-table databases, we presented in this paper an approach to computing exact consistent answers of analytic queries over star schemas under the assumption that the selection condition is independent. Our main contribution is to propose a framework for star schemas in which missing values are allowed, and in which computing the consistent answer to a usual projection-selection-join query, or to an analytic query, is quasi-linear in the size of the data warehouse, i.e. in $\mathcal{O}(W\cdot\log(W))$ where $W$ is the size of the data warehouse.
Our current work follows two main lines of research. First, the implementation of our approach is an important issue under consideration. We notice in this respect that, as m_Chase(T)m\_Chase(T) has to be ‘re-computed’ after each update, the question of incrementally implementing the changes in m_Chase(T)m\_Chase(T) is crucial regarding efficiency. This issue is currently investigated based on Algorithm 2 and on the statements in Proposition 2.
@@
Acknowledments
The authors thank the reviewers for their thorough reading of our manuscript. Their comments and suggestions helped us improve the contents and the presentation of the modified version.
@@ Appendix A Proof of Proposition 2
Proposition 2  Let $T$ be a star-table over universe $U$. The following hold:
  1. 1.
    For every $\sigma$ in m_Chase(T)m\_Chase(T) and every $i=1,\ldots,n$, if $K_{i}\in sch(\sigma)$ then $|{\sf tuples}(\sigma(K_{i}))|=1$. Consequently, if $\mathbb{K}\subseteq sch(\sigma)$ then $|{\sf tuples}(\sigma(\mathbb{K}))|=1$.
  2. 2.
    For every tuple $k$ over $\mathbb{K}$ in $\mathcal{T}$, there exists at most one $\sigma$ in m_Chase(T)m\_Chase(T) such that $\mathbb{K}\subseteq sch(\sigma)$ and ${\sf tuples}(\sigma(\mathbb{K}))=(k)$.
  3. 3.
    Moreover, m_Chase(T)m\_Chase(T) contains the following two kinds of m-tuples:
    1. (a)
      $\sigma$ for which there exists $i_{0}\in\{1,\ldots,n\}$ such that:
      • $sch(\sigma)\subseteq sch(D_{i_{0}})$, ${\sf tuples}(\sigma(K_{i_{0}}))=(k_{i_{0}})$ and for every $t\in F$, $t.K_{i_{0}}\neq k_{i_{0}}$,
      • for every $A\in sch^{*}(D_{i_{0}})$, $\sigma(A)=\{a~|~(\exists q\in D_{i_{0}})(q.K_{i_{0}}=k_{i_{0}}\wedge q.A=a)\}$.
    2. (b)
      $\sigma$ such that $\mathbb{K}\subseteq sch(\sigma)$ and ${\sf tuples}(\sigma(\mathbb{K}))=(k)$, and
      • for every $M_{j}\in\mathbb{M}$, $\sigma(M_{j})=\{m_{j}~|~(\exists t\in F)(t.\mathbb{K}=\sigma(\mathbb{K})\wedge t.M_{j}=m_{j})\}$,
      • for every $i=1,\ldots,n$, for every $A\in sch^{*}(D_{i})$,
        $\sigma(A)=\{a~|~(\exists t\in D_{i})(t.K_{i}=k.K_{i}\wedge t.A=a)\}$.
Proof. 1. The result comes from the fact that, to generate multi-valued components over an attribute $A$, the algorithm m_Chasem\_Chase considers a functional dependency whose right hand-side is $A$. As in the case of a star-table, no dependency in $FD$ has its right hand-side in $\mathbb{K}$, it impossible to generate conflicts on key attributes. The proof of this item is therefore complete.
2. Let $\sigma_{1}$ and $\sigma_{2}$ be in m_Chase(T)m\_Chase(T) such that $\mathbb{K}\subseteq sch(\sigma_{i})$ for $i=1,2$ and $\sigma_{1}(\mathbb{K})=\sigma_{2}(\mathbb{K})$. Since $\sigma_{1}\neq\sigma_{2}$ (the algorithm eliminates duplicates) there exists $A$ not in $\mathbb{K}$ such that $\sigma_{1}(A)\neq\sigma_{2}(A)$. Since $A$ is not in $\mathbb{K}$, either there is $i_{0}$ in $\{1,\ldots,n\}$ such that $A\in sch^{*}(D_{i_{0}})$ or $A$ is in $\mathbb{M}$. Hence, $FD$ contains a dependency $X\to A$ where $X=K_{i_{0}}$ in the former case, or $X=\mathbb{K}$ in the latter case. Since, as stated above, $|\sigma_{1}(X)|=|\sigma_{2}(X)|=1$, we have $\sigma_{1}(X)=\sigma_{2}(X)$. Thus, applying the m_Chase rule to m_Chase(T)m\_Chase(T) changes $\sigma_{1}(A)$ and $\sigma_{2}(A)$ into $\sigma_{1}(A)\cup\sigma_{2}(A)$. As by definition of m_Chase(T)m\_Chase(T), m_Chase(m_Chase(T))=m_Chase(T)m\_Chase(m\_Chase(T))=m\_Chase(T), we obtain a contradiction which completes the proof of this item.
3. To prove the results, we let $\left(\Sigma_{l}\right)_{l\geq 0}$ be the sequence of m-tables successively generated when running Algorithm 1, and we let $\lambda$ be the least integer such that $\Sigma_{\lambda}=\Sigma_{\lambda+1}=m\_Chase(T)$. Since $\Sigma_{0}$ is obtained from $T$ by ‘converting’ tuples in the data warehouse into m-tuples, according to the restrictions on missing values earlier stated, every m-tuple $\sigma$ in $\Sigma_{0}$ satisfies the following:
  • (A)
    If $\mathbb{K}\not\subseteq sch(\sigma)$ then $\sigma$ has been obtained from a tuple in a dimension table. Thus, there exists $i_{0}\in\{1,\ldots,n\}$ such that $sch(\sigma)\subseteq sch(D_{i_{0}})$, ${\sf tuples}(\sigma(K_{i_{0}}))=(k_{i_{0}})$ and or every $A\in sch^{*}(D_{i_{0}})$, $\sigma(A)\subseteq\{a~|~(\exists q\in D_{i_{0}})(q.K_{i_{0}}=k_{i_{0}}\wedge q.A=a)\}$.
  • (B)
    Otherwise, $\sigma$ has been obtained from a tuple in the fact table, and thus, for every $M_{j}\in\mathbb{M}$, $\sigma(M_{j})\subseteq\{m_{j}~|~(\exists t\in F)(t.\mathbb{K}=\sigma(\mathbb{K})\wedge t.M_{j}=m_{j})\}$. Moreover, for every $i=1,\ldots,n$, for every $A\in sch^{*}(D_{i})$, we have $\sigma(A)=\emptyset$ (because $A\not\in sch(F)$), implying that $\sigma(A)\subseteq\{a~|~(\exists t\in D_{i})(t.K_{i}=k.K_{i}\wedge t.A=a)\}$.
We show by induction that, for $l\geq 0$, the m-tuples in $\Sigma_{l}$ satisfy one of the two items (A) or (B) above. Indeed, assuming that all m-tuples in $\Sigma_{l}$ satisfy (A) and (B), let $\sigma$ and $\sigma^{\prime}$ be in $\Sigma_{l}$ and $Y\to B$ the dependency in $FD$ used by the m_Chase rule when generating $\Sigma_{l+1}$. Thus, $YB\subseteq sch(\sigma)$ and $YB\subseteq sch(\sigma^{\prime})$, and applying the m_Chase rule to $\sigma$ and $\sigma^{\prime}$ modifies $\sigma(B)$ and $\sigma^{\prime}(B)$ to produce $\sigma_{1}$ and $\sigma^{\prime}_{1}$ such that $\sigma_{1}(B)=\sigma^{\prime}_{1}(B)=\sigma(B)\cup\sigma^{\prime}(B)$. Since $B\not\in\mathbb{K}$, it follows that $sch(\sigma)\cap\mathbb{K}=sch(\sigma_{1})\cap\mathbb{K}$ and $sch(\sigma^{\prime})\cap\mathbb{K}=sch(\sigma^{\prime}_{1})\cap\mathbb{K}$, thus that $\sigma_{1}$ falls in case (A), respectively in case (B), if and only if $\sigma$ falls in case (A), respectively in case (B), and the same happens for $\sigma^{\prime}$ and $\sigma^{\prime}_{1}$. Moreover, since by our induction hypothesis, $\sigma$ and $\sigma^{\prime}$ satisfy the inclusions stated in items (A) and (B), it is easy to see that such is also the case for $\sigma_{1}$ and $\sigma^{\prime}_{1}$. Therefore, it turns out that the m-tuples in $\Sigma_{l+1}$ satisfy the inclusions stated in items (A) and (B) above; in particular, such is the case for m_Chase(T)m\_Chase(T).
We now prove that, when considering m_Chase(T)m\_Chase(T), the inclusions involving $\sigma(A)$ in item (A) or (B) are in fact equalities. In case of item (A), let $\sigma\in m\_Chase(T)$ and $i_{0}\in\{1,\ldots,n\}$ such that $K_{i_{0}}\in sch(\sigma)$, $\mathbb{K}\not\subseteq sch(\sigma)$. If there exists $A\in sch^{*}(D_{i_{0}})$ and $q\in D_{i_{0}}$ such that $A\in sch(q)$, $\sigma(K_{i_{0}})=q.K_{i_{0}}$ and $q.A\not\in\sigma(A)$, then by construction of m_Chase(T)m\_Chase(T) there must exist $\sigma_{q}\in m\_Chase(T)$ such that $q\in{\sf tuples}(\sigma_{q}(sch(q)))$. It follows that $\sigma(A)\neq\sigma_{q}(A)$ whereas $\sigma(K_{i_{0}})=\sigma_{q}(K_{i_{0}})$, thus that applying the m_Chase rule to $\sigma$, $\sigma_{q}$ and $K_{i_{0}}\to A$ in $FD$, changes m_Chase(T)m\_Chase(T). In other words this implies that $m\_Chase(m\_Chase(T))\neq m\_Chase(T)$, which is a contradiction. Since similar arguments hold in the case where $\sigma$ satisfies item (B), this part of the proof is complete.
The last step in this proof consists in showing that for every $\sigma$ in m_Chase(T)m\_Chase(T) of kind (a) in the proposition and such that $\sigma(K_{i_{0}})=k_{i_{0}}$, $k_{i_{0}}$ does not occur in $F$. Indeed, if such is the case, m_Chase(T)m\_Chase(T) contains $\sigma^{\prime}$ such that $\mathbb{K}\subseteq sch(\sigma^{\prime})$ and $\sigma^{\prime}(K_{i_{0}})=\sigma(K_{i_{0}})=k_{i_{0}}$. Since m_Chase(T)m\_Chase(T) is closed under the m_Chase rule, every attribute $A$ in $sch(\sigma)\setminus K_{i_{0}}$ belongs to $D_{i_{0}}^{*}$ and so, as $K_{i_{0}}\to A$ is in $FD$, AA is in $sch(\sigma^{\prime})$ and $\sigma(A)=\sigma^{\prime}(A)$. Hence, $\sigma\sqsubseteq\sigma^{\prime}$, and so, applying the m_Chase rule to $\sigma$ and $\sigma^{\prime}$ changes m_Chase(T)m\_Chase(T) by removing $\sigma$, a contradiction to the fact that m_Chase(T)m\_Chase(T) is closed under the m_Chase rule. The proof is therefore complete. $\Box$
Appendix B Proof of Lemma 1
Lemma 1  Let $T$ be a star-table over universe $U$, $S$ a subset of ${\sf True}(\mathcal{T})$ and $t$ a tuple in ${\sf Cons}(\mathcal{T})$. If $S\models FD$ then $S\cup\{t\}\models FD$.
Proof. Since $S\models FD$, every $\sigma$ in the m-table m_Chase(S)m\_Chase(S) is such that, for every $A$ in $sch(\sigma)$, $\sigma(A)=(a)$. Thus m_Chase(S)m\_Chase(S) can be seen as a table that we denote by $S^{*}$. Moreover, denoting by $S_{t}$ the table $S\cup\{t\}$, we have $m\_Chase(S_{t})=m\_Chase(S^{*}\cup\{t\})$. Let $S^{*}_{t}=S^{*}\cup\{t\}$, and let us consider the computation of $m\_Chase(S^{*}_{t})$.
- If $q_{1}$ and $q_{2}$ are in $S^{*}_{t}$, then either $q_{1}$ and $q_{2}$ are not defined over $A$ or, they are both defined such that $q_{1}.A=q_{2}.A$. In this case m_Chasem\_Chase does not change $S^{*}_{t}$.
- If $q_{1}\in S^{*}_{t}$, $q_{2}=t$, and $q_{1}$ and $q_{2}$ are not defined over $A$, then m_Chasem\_Chase does not change $S^{*}_{t}$.
- If $q_{1}\in S^{*}_{t}$, $q_{2}=t$, and $q_{1}$ and $q_{2}$ are defined over $A$. Since $t\in{\sf Cons}(\mathcal{T})$, it is not possible that $q_{1}.A\neq t.A$. Thus in this case again, m_Chasem\_Chase does not change $S^{*}_{t}$.
- If $q_{1}\in S^{*}_{t}$, $q_{2}=t$, $q_{1}$ is defined over $A$ and $q_{2}$ is not defined over $A$. Then m_Chasem\_Chase changes $t$ into $ta$ where $a=q_{1}.A$.
- If $q_{1}\in S^{*}_{t}$, $q_{2}=t$, $q_{1}$ is not defined over $A$ and $q_{2}$ is defined over $A$. Then m_Chasem\_Chase changes $q_{1}$ into $qa$ where $a=t.A$.
Indeed, let $q_{1}$ and $q_{2}$ be in $\Sigma$ and $X\to A$ be in $FD$. The only possibility for m_Chasem\_Chase to change $\Sigma$ is that $q_{1}.X=q_{2}.X$, $q_{1}.A=a$ and $q_{2}$ is not defined over $A$. As this does not happen in $S^{*}$ (because $m\_Chase(S^{*})=S^{*}$), and as this does not happen in $\Sigma$ (by definition of $\Sigma$), we obtain that $m\_Chase(\Sigma)=\Sigma$. It follows that $m\_Chase(S^{*}_{t})=\Sigma$, and since we have seen that no conflicts occur in $\Sigma$, $S^{*}_{t}\models FD$ holds. The proof is therefore complete. $\Box$
Appendix C Proof of Proposition 6
Proposition 6  Let $T$ be a star-table over universe $U$. For a given $X\to A$ in $FD$ and a given tuple $x$ over $X$, let $\{xa_{1},\ldots,xa_{q}\}$ be the set of all tuples over $XA$ that belong to ${\sf True}(\mathcal{T})$. Then, for every $R$ in ${\sf Rep}(T)$, there exists $q_{0}$ in $\{1,\ldots,q\}$ such that $xa_{q_{0}}$ is in ${\sf True}(\mathcal{R})$.
Proof. It is easy to show that the proposition holds for $q=1$. Indeed, in this case $xa_{1}$ is in ${\sf Cons}(\mathcal{T})$, and thus in ${\sf True}(\mathcal{R})$ for every $R$ in ${\sf Rep}(T)$. We now assume that $q>1$, which implies that for every $\sigma$ in m_Chase(T)m\_Chase(T) such that $XA\subseteq sch(\sigma)$ and $x$ is in ${\sf tuples}(\sigma(X))$, ${\sf tuples}(\sigma(XA))=(xa_{1},\ldots,xa_{q})$ with $q>1$. Moreover, given the form of functional dependencies in $FD$, either $(a)$ $X=\mathbb{K}$ and $A$ is in $\mathbb{M}$, or $(b)$ there exists $i_{0}$ in $\{1,\ldots,n\}$ such that $X=K_{i_{0}}$ and $A$ is in $sch^{*}(D_{i_{0}})$.
Let $R$ be in ${\sf Rep}(T)$ such that ${\sf True}(\mathcal{R})$ contains no $xa$ from $\{xa_{1},\ldots,xa_{q}\}$, and let $q_{0}$ in $\{1,\ldots,q\}$. Denoting by $R^{\prime}$ the table $R\cup\{xa_{q_{0}}\}$, we show that in either case $(a)$ or $(b)$ above, $R^{\prime}\models FD$. To this end, we first notice that, since for every $\mathbb{K}$-value $k$ and for any $K_{i}$-value $_{i}, kk$ and $k_{i}$ are in ${\sf Cons}(\mathcal{T})$, the tuple $x$ is in ${\sf True}(\mathcal{R})$. Hence, ${\sf True}(\mathcal{R})$ contains tuples whose $X$-value is $x$ and for every such tuple, the $A$-value is missing. The cases $(a)$ and $(b)$ are then handled as follows when computing $m\_Chase(R^{\prime})$: every $t$ in ${\sf True}(\mathcal{R}^{\prime})$ such that $t.X=x$ is considered and its $A$-value is set to $a_{q_{0}}$. Denoting by $\Sigma$ the resulting m-table, let $X^{\prime}\to A^{\prime}$ be a dependency in $FD$, and let $\sigma_{1}$ and $\sigma_{2}$ be two m-tuples in $\Sigma$ such that $\sigma_{1}(X^{\prime})\cap\sigma_{2}(X^{\prime})\neq\emptyset$ where $X^{\prime}$ is either $\mathbb{K}$ or $K_{j}$. Since $R$ is in ${\sf Rep}(T)$, $R$ satisfies $FD$ and thus, all components of all m-tuples in m_Chase(T)m\_Chase(T) are singletons. Since this also holds in $\Sigma$, it follows that $\sigma_{1}(X^{\prime})\cap\sigma_{2}(X^{\prime})\neq\emptyset$ implies that $\sigma_{1}(X^{\prime})=\sigma_{2}(X^{\prime})$. Then, the following holds:
- If $X^{\prime}=X$ then by definition of $\Sigma$, $\sigma_{1}(A)=\sigma_{2}(A)=(a_{q_{0}})$. Thus $\Sigma$ is not changed.
- If $X^{\prime}=K_{j}$ and $X=\mathbb{K}$, since as above we have $A^{\prime}\neq A$, m_Chasem\_Chase does not change $\Sigma$.
- If $X^{\prime}=\mathbb{K}$ and $X=K_{i}$, then in ${\sf True}(\mathcal{R}^{\prime})$, we have $\sigma_{1}(X)=\sigma_{2}(X)=x$, and $A$ is neither in $sch(\sigma_{1})$ nor in $sch(\sigma_{2})$. By definition of $\Sigma$, $\sigma_{1}(A)$ and $\sigma_{2}(A)$ are set to $(a_{q_{0}})$, and then here again, m_Chasem\_Chase does not change $\Sigma$.
We thus obtain that $\Sigma=m\_Chase(\mathcal{R}^{\prime})$ and that $R^{\prime}\models FD$, because $R\models FD$ and no conflicts occur in $\Sigma$. Moreover, as ${\sf True}(\mathcal{R})\subseteq{\sf True}(\mathcal{R}^{\prime})$, we have ${\sf Cons}(\mathcal{T})\subseteq{\sf True}(\mathcal{R}^{\prime})$. It follows from Definition 2 that $R$ can not be in ${\sf Rep}(T)$ since the inclusion ${\sf True}(\mathcal{R})\subset{\sf True}(\mathcal{R}^{\prime})$ is strict (because $xa_{q_{0}}$ is in ${\sf True}(\mathcal{R}^{\prime})$ but not in ${\sf True}(\mathcal{R})$). As this is a contradiction with our hypothesis that $R$ is in ${\sf Rep}(T)$, the proof is complete. $\Box$
Appendix D Proof of Proposition 7
@@
Indeed, let $X\to A$ in $FD$, $t_{\varphi}(\sigma)$ and $t_{\varphi}(\sigma^{\prime})$ in $T_{\varphi}$ such that $t_{\varphi}(\sigma).X=t_{\varphi}(\sigma^{\prime}).X$. If $\sigma$ and $\sigma^{\prime}$ are defined over $A$, then, by construction of $T_{\varphi}$, we have $t_{\varphi}(\sigma).A=t_{\varphi}(\sigma^{\prime}).A$, in which case m_Chasem\_Chase does not change $T_{\varphi}$. The case whereby $\sigma$ is defined over $A$ and $\sigma^{\prime}$ is not defined over $A$ is not possible in m_Chase(T)m\_Chase(T), and thus it is not possible that $t_{\varphi}(\sigma)$ is defined over $A$ while $t_{\varphi}(\sigma^{\prime})$ is not. Therefore, we have $m\_Chase(T_{\varphi})=T_{\varphi}$, and this implies that $T_{\varphi}\models FD$, because no conflicts can occur in $T_{\varphi}$.
@@
Recalling from LauS25 that, when $FD$ is acyclic, we also have $m\_Chase({\sf Cons}(\mathcal{T}))={\sf Cons}(\mathcal{T})$ and ${\sf Cons}(\mathcal{T})\models FD$, we prove by induction that, at each step $k$ of the computation of $m\_Chase(R_{\varphi})$ the obtained m-table $\Sigma^{k}$ is such that $(1)$ $\Sigma^{k}$ contains no conflict and $(2)$ $\tau(\Sigma^{k})=\tau(R_{\varphi})$.
@@
- If $q$ and $q^{\prime}$ are both in $T_{\varphi}$ (respectively both in ${\sf Cons}(\mathcal{T})$): as these sets are closed under the m_Chase rule, $t$ and $t^{\prime}$ are both in $T_{\varphi}$ (respectively both in ${\sf Cons}(\mathcal{T})$), and so $t^{\prime}_{A}$ is in $T_{\varphi}$ (respectively in ${\sf Cons}(\mathcal{T})$), because $t^{\prime}_{A}$ is obtained through the m_Chase rule.
- If $q$ is in ${\sf Cons}(\mathcal{T})$ and $q^{\prime}$ is in $T_{\varphi}\cap{\sf Confl}(\mathcal{T})$: in this case $xa$ is in ${\sf Cons}(\mathcal{T})$ (because $xa\sqsubseteq q$). If $\sigma^{\prime}$ is the m-tuple in m_Chase(T)m\_Chase(T) such that $q^{\prime}=t_{\varphi}(\sigma^{\prime})$, $XA\subseteq sch(\sigma^{\prime})$ and $xa$ is in ${\sf tuples}(\sigma^{\prime}(XA))$. Hence $t^{\prime}_{A}\sqsubseteq t_{\varphi}(\sigma^{\prime})$, and thus $t^{\prime}_{A}$ is in ${\sf tuples}(R_{\varphi})$.
- If $q$ is in $T_{\varphi}\cap{\sf Confl}(\mathcal{T})$ and $q^{\prime}$ in ${\sf Cons}(\mathcal{T})$: in this case, assume first that $xa$ is in ${\sf Cons}(\mathcal{T})$. Then, $t^{\prime}_{A}$ is in ${\sf Cons}(\mathcal{T})$, because for every $Y\to B$ in $FD$ other than $X\to A$, if $YB\subseteq sch(t^{\prime}_{A})$ then $YB\subseteq sch(t^{\prime})$.
Assuming now that $xa$ is in ${\sf Confl}(\mathcal{T})$, we have $a=\varphi(x)$ and so, for every $Y\to B$ such that $YB\subseteq sch(t^{\prime}_{A})$, $t^{\prime}_{A}.B=\varphi(t^{\prime}_{A}.Y)$ (because as $t^{\prime}$ is in ${\sf Cons}(\mathcal{T})$, for every $Y\to B$ such that $YB\subseteq sch(t^{\prime})$, $t^{\prime}.B=\varphi(t^{\prime}.Y)$). Hence, there exists $\sigma^{\prime}$ in m_Chase(T)m\_Chase(T) such that $t^{\prime}_{A}\sqsubseteq t_{\varphi}(\sigma^{\prime})$. Therefore $t^{\prime}_{A}$ is in $\tau(R_{\varphi})$.
@@
To this end, we let $S$ be such that ${\sf True}(\mathcal{R}_{\varphi})\subseteq{\sf True}(\mathcal{S})\subseteq{\sf True}(\mathcal{T})$ and $S\models FD$, and we let $q$ be a tuple in ${\sf True}(\mathcal{S})\setminus{\sf True}(\mathcal{R}_{\varphi})$. As $q$ is in ${\sf True}(\mathcal{T})$, there exist $\sigma$ in m_Chase(T)m\_Chase(T) and $t$ in ${\sf tuples}(\sigma)$ such that $q\sqsubseteq t$. Since ${\sf Cons}(\mathcal{T})\subseteq{\sf True}(\mathcal{R}_{\varphi})$, it follows that $q$ is in ${\sf Confl}(\mathcal{T})$, implying that there exists $X\to A$ in $FD$ such that $XA\subseteq sch(q)$ and ${\sf True}(\mathcal{T})$ contains $xa^{\prime}$ such that $x=q.X$ and $a^{\prime}\neq q.A$. By construction of $R_{\varphi}$, ${\sf True}(\mathcal{R}_{\varphi})$ contains a tuple $t^{\prime}$ from ${\sf tuples}(\sigma)$, and so, we have $t.X=q.X=t^{\prime}.X$ (because for every $X\to A\in FD$, $|{\sf tuples}(\sigma(X))|=1$) and $t.A\neq t^{\prime}.A$. As $t.A=q.A$ and $t^{\prime}$ is in ${\sf True}(\mathcal{S})$, this implies that $S\not\models FD$, which is a contradiction. This part of the proof is therefore complete.
@@
Appendix E Proof of Proposition 9
Proposition 9  Let $T$ be a star-table over universe $U$, and let $Q:{\tt select}$ $X$ from $T$ where $\Gamma$ be a query such that $\Gamma$ is an independent selection condition. Then ${\sf C\_ans}(Q)$ is the set of all tuples $x$ over $X$ for which there exists $\sigma$ in m_Chase(T)m\_Chase(T) such that